# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 12 Conclusions and outlook
We have reviewed the statistical properties of AH and NAH theories, describing multicomponent scalar fields coupled with Abelian and non-Abelian gauge fields, mostly in three dimensions. As remarked several times, the interplay between local gauge and global symmetries determines the nature of the Higgs phases and of the phase transitions of these lattice gauge systems. Their nature also crucially depends on the role played by gauge modes at criticality, whether they are critical at the transition or not. In the latter case, their only role is that of restricting the critical scalar modes to the gauge-invariant sector. The topological properties of the gauge modes can also give rise to topological transitions between Higgs phases with the same scalar global symmetry. These transitions are related to the breaking pattern of the gauge symmetry in the Higgs phase. We discuss examples with a residual ${\mathbb{Z}}_{Q}$ gauge symmetry, leading to topological ${\mathbb{Z}}_{Q}$ gauge transitions when varying the gauge coupling.
@@
The results reviewed here show that the possible critical behaviors observed in 3D lattice gauge Higgs systems are quite varied, showing features that cannot be explained by the standard LGW paradigm. All continuous transitions observed in LAH and LNAH systems and discussed in this review can be classified as follows, as outlined in detail in Sec. 3.1: (i) LGW transitions with gauge-invariant order parameter and noncritical gauge modes; (ii) LGW× transitions, where gauge modes are also not critical, but where the order-parameter field is gauge dependent; (iii) GFT transitions, where gauge modes are critical and that require an effective description that includes the gauge fields; (iv) topological transitions driven only by topological gauge modes, without any local gauge-invariant scalar order parameter.
@@
We have focused on classical statistical systems, whose partition function is defined by integrating over lattice configurations. However, the presented results in ($d+1$) dimensions also apply to some quantum transitions in $d$ dimensions, using the quantum-to-classical mapping. In particular, they are relevant for $d$-dimensional quantum transitions, which can be related with transitions in $(d+1)$-dimensional classical isotropic systems. In particular, this requires that their dynamic exponent satisfies $z=1$. However, further interesting developments may come from a thorough analysis of quantum transitions in the presence of emerging gauge symmetries with $z\neq 1$ [14, 101]. This issue definitely calls for further studies.
@@
We finally mention that new scenarios may arise from the inclusion of fermionic fields in AH and NAH systems. This topic is relevant in condensed-matter physics as many field-theoretical models with emerging gauge symmetries include both fermionic and scalar excitations, see, e.g., Refs. [96, 20, 452, 453, 235, 372]. The study of the corresponding lattice gauge systems may extend the phenomenology of phase transitions in the presence of gauge symmetries. As for LAH and LNAH systems with only scalar matter, one key issue would be that of identifying the 3D critical behaviors realizing the continuous limits of the corresponding SFTs. The same issue is also important in lattice gauge models with fermionic matter only. Some studies of the phase diagram and critical behavior of 3D models including fermionic fields can be found in Refs. [454, 455, 456, 457, 458, 459, 460, 461, 462, 463, 464, 76, 465, 466, 467, 468, 469, 470, 471, 472, 473, 474, 475, 476].
@@ Appendix A Renormalization-group theory of critical phenomena
In this appendix we report an overview of the RG predictions for the scaling behavior at continuous phase transitions [129, 94, 113, 114, 124, 125, 130, 6, 132, 133, 134, 135, 136, 127, 2, 8, 9, 137, 11, 95], which have been thoroughly checked in various analytic and numerical studies. We present the RG scaling relations in the thermodynamic (infinite-volume) limit and in the FSS limit and, in particular, we discuss the leading scaling behavior and the scaling corrections characterizing the deviations from the asymptotic relations. We only consider classical transitions, generally driven by thermal fluctuations; the extension to quantum transitions driven by quantum fluctuations is discussed in Refs. [118, 14, 138, 101]. FSS behaviors also emerge at first-order transitions. These phenomena for classical and quantum transitions are discussed in Refs. [117, 477, 115, 478]. In the following we do not discuss the critical dynamics at thermal and quantum transitions. Reviews can be found in Refs. [479, 480, 101].
@@ A.1 Universal power laws at the critical point
To fix the ideas, we consider a prototypical $d$-dimensional statistical system undergoing a continuous transition. The behavior of the system is controlled by two relevant Hamiltonian parameters $r$ and $h$, which can be defined so as to vanish at the critical point. Thus, the critical point is located at $r=h=0$. We also assume the presence of a ${\mathbb{Z}}_{2}$-symmetry, as it occurs, e.g., in Ising transitions, which is unbroken in the paramagnetic phase with $r>0$ and spontaneously broken in the ferromagnetic phase with $r<0$. The parameter $r$ is associated with a RG perturbation that is invariant under the global symmetry and is usually identified with the reduced temperature $r\sim T/T_{c}-1$. The parameter $h$ is assumed to be odd with respect to the ${\mathbb{Z}}_{2}$-symmetry, thus it can be identified with an external homogeneous field coupled with the order parameter driving the symmetry breaking.
In the thermodynamic limit, when approaching the critical point from the disordered phase keeping $h=0$, the length scale $\xi$ of the critical modes diverges as $\xi\sim r^{-\nu}$, where $\nu$ is a universal length-scale critical exponent. Another universal critical exponent $\eta$ is traditionally introduced to characterize the space dependence at criticality of the two-point function of the order parameter: $G(\bm{x}_{1},\bm{x}_{2})\sim|\bm{x_{1}}-\bm{x_{2}}|^{-(d-2+\eta)}$ for $r=h=0$. The RG dimensions of the perturbations associated with $r$ and $h$ are related to the critical exponents $\nu$ and $\eta$, as [125, 95]
@@
Using scaling and hyperscaling arguments (see, e.g., Refs. [114, 2, 95]), the exponents associated with the scaling behavior of other observables, like the magnetization or the critical equation of state, can be expressed in terms of the two independent exponents $\nu$ and $\eta$. The corrections to the asymptotic power laws are generally suppressed as $\xi^{-\omega}$, where $\omega>0$ is another universal exponent characterizing the critical behavior [2, 95].
@@ A.2 Scaling laws in the thermodynamic limit
where the summation is over all configurations $\{\varphi\}$ of the system. According to the RG theory, $F$ obeys a general scaling law. Indeed, we can write the $F$ in terms of the nonlinear scaling fields associated with the RG eigenoperators at the fixed point of the RG flow [125, 95]. Therefore, close to a continuous transition, the free-energy density in the thermodynamic infinite-volume limit can be written in terms of scaling fields [125], as
@@
Here $F_{\rm reg}$ is a nonuniversal function, which is analytic at the critical point and must be even with respect to the parameter $h$ related to the odd perturbation (for a general symmetry group it should depend on a combination of $h$ that is invariant under the symmetry transformations). The singular contribution, $F_{\rm sing}$, bears the nonanalyticity of the critical behavior and its universal features. The arguments of $F_{\rm sing}$ are the so-called nonlinear scaling fields [125]: they are analytic nonlinear functions of the model parameters, related to the eigenoperators that diagonalize the RG flow close to the RG fixed point, and transform multiplicatively under the RG flow. We have singled out the relevant scaling fields $u_{r}$ and $u_{h}$, that are associated with the model parameters $r$ and $h$ and have positive RG dimensions $y_{r}$ amd $y_{h}$. The scaling fields $\{v_{i}\}$ (there is an infinite number of them) are RG irrelevant, i.e., their RG dimensions $y_{i}$ are negative. They give rise to scaling corrections to the asymptotic critical behavior in the infinite-volume limit. Assuming that they are ordered so that $|y_{1}|\leq|y_{2}|,\leq\ldots$, the RG dimension $y_{1}<0$ of the leading irrelevant perturbation generally determines the leading scaling corrections and we identify $\omega=-y_{1}$.
In general, the nonlinear scaling fields depend on the control parameters $r$ and $h$. Taking into account the assumed ${\mathbb{Z}}_{2}$ symmetry and the respectively even and odd properties of $r$ and $h$, close to the critical point the relevant scaling fields $u_{r}$ and $u_{h}$ can be generally expanded as
@@
where $c_{r}$ and $c_{h}$ are nonuniversal constants. As for the irrelevant scaling fields $v_{i}$, they are usually nonvanishing at the critical point.
@@
where ${\cal F}_{0}$ and ${\cal F}_{\omega}$ are universal scaling functions, and we only kept only the leading correction to scaling, decaying the exponent $y_{1}=-\omega$. Note, however, that, since the scaling fields have arbitrary normalization, the universality of the scaling functions ${\cal F}_{0}$ and ${\cal F}_{\omega}$ holds apart from a normalization of each argument and an overall constant. By differentiating the free energy density, one can straightforwardly derive analogous scaling relations for other observables, such as the energy density and the magnetization. To eventually obtain the scaling relations in terms of the external parameters $r$ and $h$ controlling the critical behavior, one needs to expand the scaling fields as in Eq. (135). The subleading terms in these expansions give rise to analytic scaling corrections.
Analogous scaling relations can be obtained for the correlation functions of local operators. For example, the two-point correlation function of the local operator $O({\bm{x}})$, whose RG dimension is $y_{o}$, is expected to asymptotically behave as
@@ A.3 Finite-size scaling
Finite-size effects in critical phenomena have been the object of many theoretical studies [481, 482, 135, 136, 134, 95]. The FSS theory describes the critical behavior around a critical point, when the correlation length $\xi$ of the critical modes becomes comparable with the size $L$ of the system. The FSS approach is one of the most effective techniques to determine the universal features of continuous phase transitions, for instance, the critical exponents. While infinite-volume methods require the condition $\xi\ll L$ to be satisfied, FSS methods apply in the less demanding regime $\xi\sim L$. The FSS theory provides the asymptotic scaling behavior when both $L,\xi\to\infty$, keeping their ratio $\xi/L$ fixed. This regime presents universal features, shared by all systems whose transition belongs to the same universality class.
A.3.1 Finite-size scaling of the free-energy density
In the FSS framework, the finite size of the system is taken into account by adding the length scale $L$ in the scaling laws [125, 483, 135, 136, 118, 484, 95, 138, 101]. In the following we focus for concreteness on finite-size systems without boundaries, such as systems with periodic or antiperiodic boundary conditions. In these cases the scaling laws, and in particular the corrections to the asymptotic scaling behaviors, are simpler. The corrections to the bulk scaling laws due to the presence of boundaries are discussed in, e.g., Refs. [135, 136, 484, 138, 101]. In the FSS limit the free-energy density can be written
@@ A.3.1 Finite-size scaling of the free-energy density
where $F_{\rm reg}$ is again a nonuniversal analytic function, which is assumed to be independent of $L$, or, more plausibly, to depend on $L$ only through exponentially small corrections [135, 136, 484, 138]. According to the RG theory, the singular part $F_{\rm sing}$ satisfies the homegeneous scaling law
@@
The above scaling laws hold in generic systems. However, in some cases the scaling behavior is more complex, due to the appearance of logarithmic terms [125]. They may be induced by the presence of marginal RG perturbations, as it happens in BKT transitions in 2D U(1)-symmetric systems [485, 413, 486, 444, 445], or by resonances between the RG eigenvalues, as it occurs in 2D Ising transitions [125, 487], or in 3D O$(N)$-vector models in the large-$N$ limit [488, 489, 490].
@@
We finally remark that, in numerical and experimental investigations of finite-size systems, the knowledge of the leading asymptotic behavior may not be enough to estimate the critical parameters, because data are generally available for limited ranges of parameter values and system sizes, which are often relatively small. Under these circumstances, the asymptotic FSS predictions are affected by sizable scaling corrections. Accurate estimates of the critical parameters thus require a robust control of the corrections to the asymptotic scaling behavior. This is also important for a conclusive identification of the universality class of a continuous transition, when there are no solid theoretical arguments to predict it. Moreover, an understanding of the finite-size effects is relevant for the experiments, when relatively small systems are considered (see, e.g., Ref. [491]), or in particle systems trapped by external (usually harmonic) forces, as in cold-atom setups (see, e.g., Refs. [492, 493, 494, 495, 496, 497]).
A.3.2 Finite-size scaling of the order-parameter correlation function

### v3

@@ 12 Conclusions and outlook
We have reviewed the statistical properties of lattice Abelian and non-Abelian Higgs theories, in which multicomponent scalar fields are coupled with Abelian and non-Abelian gauge fields, mostly in three dimensions. As remarked several times, the interplay between local and global symmetries determines the nature of the Higgs phases and of the phase transitions of these lattice gauge systems. Their nature also crucially depends on the role played by gauge modes at criticality, whether they are critical at the transition or not. In the latter case, their role is that of selecting the critical scalar modes, so that only critical correlations associated with gauge-invariant scalar modes can be observed. The topological properties of the gauge modes can also give rise to topological transitions between Higgs phases with the same scalar global symmetry. These transitions are related to the breaking pattern of the gauge symmetry in the Higgs phase. We discuss examples with a residual ${\mathbb{Z}}_{Q}$ gauge symmetry, leading to topological ${\mathbb{Z}}_{Q}$ gauge transitions when varying the gauge coupling.
@@
The results reviewed here show that the possible critical behaviors observed in 3D lattice gauge Higgs systems are quite varied, showing also features that cannot be explained by the standard LGW paradigm. The continuous transitions of the statistical Abelian and non-Abelian gauge systems considered in this review can be classified as follows (see Sec. 3.1): (i) LGW transitions with gauge-invariant order parameter and noncritical gauge modes; (ii) LGW× transitions, where gauge modes are also not critical, but where the effective order-parameter field is gauge dependent; (iii) GFT transitions, where gauge modes are critical and that require an effective description that includes the gauge fields; (iv) topological transitions driven only by topological gauge modes, without any local gauge-invariant scalar order parameter.
@@
We have focused on the equilibrium thermodynamic behavior of classical statistical systems, whose partition function is defined as a path-integral functional, by integrating over continuum or lattice configurations. However, the results for classical statistical systems in ($d+1$) dimensions also apply to quantum transitions in $d$ dimensions, using the quantum-to-classical mapping. Indeed, they are relevant for the $d$-dimensional quantum transitions that can be related with classical thermal transitions in $(d+1)$-dimensional isotropic systems. In particular, this requires that the critical exponent $z$, associated with the vanishing gap $\Delta\sim\xi^{-z}$ at the quantum critical point, satisfies $z=1$. However, further interesting developments may come from a thorough analysis of quantum transitions in the presence of emerging gauge symmetries with $z\neq 1$ [14, 102]. This issue definitely calls for further studies.
Another interesting issue is the critical dynamics of statistical systems with gauge symmetries. On one side, one would like to understand if the presence of gauge symmetries gives rise to distinctive new features for the different types of dynamics that are usually considered in standard statistical systems [502, 115, 503]. A systematic study of the critical dynamics in the presence of gauge symmetries may provide further interesting characterizations of the different types of transitions that we have outlined in Sec. 3.1. On the other side, it would be interesting to understand whether it is possible to define dynamics that are not considered in the standard classification of systems without gauge symmetries. As an example, one may consider the purely relaxational dynamics with only dissipative couplings and no conservation laws, usually called dynamic model A [502]. This dynamics can be realized by considering a relaxational stochastic Langevin dynamics, which is also at the basis of the stochastic quantization of gauge theories [504, 505, 506, 507, 508, 10]. It would be interesting to understand how the expected critical slowing down, signalled by the critical divergence of the relaxation time $\tau$ of the critical modes (at criticality, $\tau\sim\xi^{z}$, where $z$ is a dynamic critical exponent), depends on type of transition, and how the nonperturbative results compare with the field-theoretical predictions that are typically obtained in the presence of a gauge fixing. Some studies of the critical dynamics have been already reported, see, e.g., Refs. [273, 509, 510, 511, 512, 513, 514, 515, 516, 277, 517, 518], most of them addressing the dynamics at superconducting phase transitions, and the topological transition in the 3D lattice ${\mathbb{Z}}_{2}$ gauge model. However, a systematic analysis of dynamic phenomena is still missing.
From a methodological point of view, most of the results presented here have been obtained by using numerical approaches. We have presented the results of the analyses of numerical data obtained in MC simulations of lattice systems, field-theoretical perturbative determinations (in the $\varepsilon$-expansion framework) of the RG flow, and nonperturbative predictions obtained in the limit of a large number of scalar components. However, other methods may also provide interesting insights in the future. We mention, e.g., the tensor renormalization-group method (see, e.g., Ref. [519] and references therein), and the bootstrap method [520, 94], which has been mostly applied to systems without gauge symmetries, see results reported in the tables of B (some results for QED3 with fermions are reviewed in Ref. [94]). Another interesting possibility consists in the use of optical lattices and trapped ions to simulate lattice Abelian and non-Abelian gauge theories, see, e.g., Refs. [521, 522, 523, 524, 525] and the reviews, Refs. [493, 494].
@@
We finally mention that new scenarios may arise from the inclusion of fermionic fields in LAH and LNAH systems. This topic is relevant in condensed-matter physics as many field-theoretical models with emerging gauge symmetries include both fermionic and scalar excitations, see, e.g., Refs. [97, 20, 526, 527, 266, 404]. The study of the corresponding lattice gauge systems may extend the phenomenology of phase transitions in the presence of gauge symmetries. As for LAH and LNAH systems with only scalar matter, one key issue would be that of identifying the 3D transitions where both matter and gauge fields are critical, allowing one to define the continuum limit of the corresponding SFTs. The same issue is also important in lattice gauge models with fermionic matter only. Some studies of the phase diagram and critical behavior of 3D models including fermionic fields can be found in Refs. [528, 529, 530, 531, 532, 533, 534, 535, 536, 537, 538, 76, 539, 540, 541, 542, 543, 544, 545, 546, 547, 548, 549, 550, 551, 552, 553, 554].
Acknowledgements
We thank Alessio Franchi, Niccolò Francini, Ivan Soler Calero, Giacomo Bracci Testasecca for collaborating with us on some of the topics considered in this review, and Alessandro Vichi for useful discussions.
@@ Appendix A Renormalization-group theory of critical phenomena
In this appendix we report an overview of the RG theory, which provides a general framework to explain the observed scaling behavior at continuous phase transitions [130, 95, 114, 115, 125, 126, 131, 6, 133, 134, 135, 136, 137, 128, 2, 8, 9, 138, 11, 96]. We present the RG scaling relations in the thermodynamic (infinite-volume) limit and in the FSS limit and, in particular, we discuss the leading scaling behavior and the scaling corrections characterizing the deviations from the asymptotic relations. We only consider classical transitions, generally driven by thermal fluctuations; the extension to quantum transitions driven by quantum fluctuations is discussed in Refs. [119, 14, 139, 102]. FSS behaviors also emerge at first-order transitions [116, 118]. These phenomena for classical and quantum transitions are discussed in Refs. [118, 555, 116, 556]. In the following we do not discuss the critical dynamics at thermal and quantum transitions. Reviews can be found in Refs. [502, 503, 102].
@@ A.1 Universal power laws at the critical point
To fix the ideas, we consider a prototypical $d$-dimensional statistical system undergoing a continuous transition. The behavior of the system is controlled by two relevant Hamiltonian parameters $r$ and $h$, which can be defined so as to vanish at the critical point. Thus, the critical point is located at $r=h=0$. We also assume the presence of a ${\mathbb{Z}}_{2}$-symmetry, as it occurs, e.g., in Ising transitions, which is unbroken in the phase with $r>0$ (paramagnetic phase in magnetic systems) and spontaneously broken in the (ferromagnetic) phase with $r<0$. The parameter $r$ is associated with a RG perturbation that is invariant under the global symmetry and is usually identified with the reduced temperature $r\sim T/T_{c}-1$. The parameter $h$ is assumed to be odd with respect to the ${\mathbb{Z}}_{2}$-symmetry, thus it can be identified with an external homogeneous field coupled with the order parameter driving the symmetry breaking.
In the thermodynamic limit, when approaching the critical point from the disordered phase keeping $h=0$, the length scale $\xi$ of the critical modes diverges as $\xi\sim r^{-\nu}$, where $\nu$ is a universal length-scale critical exponent. Another universal critical exponent $\eta$ is traditionally introduced to characterize the space dependence at criticality of the two-point function of the order parameter: $G(\bm{x}_{1},\bm{x}_{2})\sim|\bm{x_{1}}-\bm{x_{2}}|^{-(d-2+\eta)}$ for $r=h=0$. The RG dimensions of the perturbations associated with $r$ and $h$ are related with the critical exponents $\nu$ and $\eta$, as [126, 96]
@@
Using scaling and hyperscaling arguments (see, e.g., Refs. [115, 2, 96]), the exponents associated with the scaling behavior of other observables, like the magnetization or the critical equation of state, can be expressed in terms of the two independent exponents $\nu$ and $\eta$. The corrections to the asymptotic power laws behave as $\xi^{-\omega}$ for $\xi\to\infty$, where $\omega>0$ is another universal exponent characterizing the critical behavior [2, 96], which is generally associated with the leading irrelevant RG perturbation.
@@ A.2 Scaling laws in the thermodynamic limit
where the summation is over all configurations $\{\varphi\}$ of the system. According to the RG theory, $F$ obeys a general scaling law. Indeed, we can write $F$ in terms of the nonlinear scaling fields associated with the RG eigenoperators at the FP of the RG flow [126, 96]. Therefore, close to a continuous transition, the free-energy density in the infinite-volume limit can be written in terms of scaling fields [126], as
@@
Here $F_{\rm reg}$ is a nonuniversal function, which is analytic at the critical point and must be even with respect to the parameter $h$ related with the odd perturbation (for a general symmetry group it should depend on a combination of $h$ that is invariant under the symmetry transformations). The singular contribution, $F_{\rm sing}$, bears the nonanalyticity of the critical behavior and its universal features. The arguments of $F_{\rm sing}$ are the so-called nonlinear scaling fields [126]: they are analytic nonlinear functions of the model parameters, related with the eigenoperators that diagonalize the RG flow close to the FP, and transform multiplicatively under the RG flow. We have singled out the relevant scaling fields $u_{r}$ and $u_{h}$ that are associated with the model parameters $r$ and $h$, and have positive RG dimensions $y_{r}$ amd $y_{h}$. The scaling fields $\{v_{i}\}$ (there is an infinite number of them) are RG irrelevant, i.e., their RG dimensions $y_{i}$ are negative. They give rise to scaling corrections to the asymptotic critical behavior in the infinite-volume limit. Assuming that they are ordered so that $|y_{1}|\leq|y_{2}|,\leq\ldots$, the RG dimension $y_{1}<0$ generally determines the leading scaling corrections, thus we identify $\omega=-y_{1}$.
The nonlinear scaling fields depend on the control parameters $r$ and $h$. Taking into account the assumed ${\mathbb{Z}}_{2}$ symmetry and the respectively even and odd properties of $r$ and $h$, close to the critical point the relevant scaling fields $u_{r}$ and $u_{h}$ can be generally expanded as
@@
where $c_{r}$ and $c_{h}$ are nonuniversal constants. As for the irrelevant scaling fields $v_{i}$, they are generally nonvanishing at the critical point.
@@
where ${\cal F}_{0}$ and ${\cal F}_{\omega}$ are universal scaling functions, and we kept only the leading correction to scaling, which vanishes as $u_{r}^{\omega\nu}$ with $\omega=-y_{1}$ approaching the critical point. Note that, since the scaling fields have arbitrary normalization, the universality of the scaling functions ${\cal F}_{0}$ and ${\cal F}_{\omega}$ holds apart from a normalization of each argument and an overall constant. By differentiating the free-energy density, one can straightforwardly derive analogous scaling relations for other observables, such as the energy density and the magnetization. To eventually obtain the scaling relations in terms of the external parameters $r$ and $h$ controlling the approach to the critical point, one needs to expand the scaling fields as in Eq. (138). The subleading terms in these expansions give rise to analytic scaling corrections.
Analogous scaling relations can be obtained for the correlation functions of local operators. For example, the two-point correlation function of a local operator $O({\bm{x}})$, whose RG dimension is $y_{o}$, is expected to asymptotically behave as
@@ A.3 Finite-size scaling
Finite-size effects in critical phenomena have been the object of many theoretical studies [557, 558, 136, 137, 135, 96]. The FSS theory describes the critical behavior around a critical point, when the correlation length $\xi$ of the critical modes becomes comparable with the size $L$ of the system. The FSS approach is one of the most effective techniques to determine the universal features of continuous phase transitions, such as the critical exponents. While infinite-volume methods require the condition $1\ll\xi\ll L$ to be satisfied, FSS methods apply in the less demanding regime $\xi\sim L\gg 1$. The FSS theory provides the asymptotic scaling behavior when $L,\xi\to\infty$, keeping their ratio $\xi/L$ fixed. This regime presents universal features, shared by all systems whose transition belongs to the same universality class.
A.3.1 Free-energy density
In the FSS framework, the finite size of the system is taken into account by adding the length scale $L$ in the scaling laws [126, 559, 136, 137, 119, 560, 96, 139, 102]. In the following we focus for concreteness on finite-size systems without boundaries, such as systems with periodic or antiperiodic boundary conditions. In these cases the scaling laws, and in particular the corrections to the asymptotic scaling behavior, are simpler. The corrections to the bulk scaling laws due to the presence of boundaries are discussed in, e.g., Refs. [136, 137, 560, 139, 102]. In the FSS limit the free-energy density can be written as
@@ A.3.1 Free-energy density
where $F_{\rm reg}$ is again a nonuniversal analytic function, which is assumed to be independent of $L$, or, more plausibly, to depend on $L$ only through exponentially small corrections [136, 137, 560, 139]. According to the RG theory, the singular part $F_{\rm sing}$ satisfies the homogeneous scaling law
@@
The above scaling laws hold in generic systems. However, in some cases the scaling behavior is more complex, due to the appearance of logarithmic terms [126]. They may be induced by the presence of marginal RG perturbations, as it happens for the Berezinskii-Kosterlitz-Thouless (BKT) transitions that occur in 2D U(1)-symmetric systems [561, 457, 562, 488, 489], or by resonances between the RG eigenvalues, as it occurs in 2D Ising transitions [126, 563], or in 3D O$(N)$-vector models in the large-$N$ limit [564, 565, 566].
@@
We finally remark that, in numerical and experimental investigations of finite-size systems, the knowledge of the leading asymptotic behavior may not be enough to estimate the critical parameters, because data are generally available only for limited ranges of parameter values and system sizes, which are often relatively small. Under these circumstances, the asymptotic FSS predictions may be affected by sizable scaling corrections. Accurate estimates of the critical parameters thus require a robust control of the corrections to the asymptotic scaling behavior. This is also important for a conclusive identification of the universality class of a continuous transition, when there are no solid theoretical arguments to predict it. Moreover, an understanding of the finite-size effects is relevant for experiments, when relatively small systems are considered (see, e.g., Ref. [567]), or in particle systems trapped by external (usually harmonic) forces, as in cold-atom setups (see, e.g., Refs. [568, 569, 570, 571, 572, 573]).
A.3.2 Energy cumulants
As discussed in this review, statistical models with gauge symmetries also undergo topological transitions characterized by the absence of a local order parameter, such as those belonging to the 3D ${\mathbb{Z}}_{Q}$ gauge and IXY universality classes (see Sec. 3.5), which are also found in LAH and LNAH models (see Secs. 5 and 6). The numerical study of the critical properties at these topological transitions can be performed by analyzing the scaling properties of the energy cumulants, which are always gauge invariant and well defined, see, e.g., Refs. [63, 86, 205, 92].
where $Z$ is the partition function defined in Eq. (136). [^54] The cumulants $C_{k}$ can be related to the energy central moments $M_{k}=\langle\,(H-\langle H\rangle)^{k}\,\rangle$, by $C_{1}=-\langle H\rangle/V$, $C_{2}=M_{2}/V$, $C_{3}=-L^{-3}M_{3}$, $C_{4}=(M_{4}-3M_{2}^{2})/V$, etc. Note that $C_{1}=-E$ where $E=\langle H\rangle/V$ is the energy density, and $C_{2}$ is proportional to the specific heat.
As one can easily check, the first cumulant $C_{1}=-E$ is dominated by the analytic background contribution $B_{1}$, due to the relation $\nu>1/d>0$, which is generally satisfied at continuous transitions [96]. This makes the scaling behavior of the energy density subleading, and therefore rather difficult to observe (this issue has been addressed in Refs. [574, 575]). Also the second cumulant $C_{2}$, which is proportional to the specific heat, is dominated by the analytic background when the specific-heat exponent $\alpha=2\nu-d$ is negative. The singular scaling part is always the dominant term in the higher energy cumulants $C_{k}$, i.e., for $k\geq 3$. Taking into account that their numerical computation becomes harder and harder with increasing $k$, due to significant cancellations which make their relative error quite large and increasing with $k$, optimal results are obtained by focusing on $C_{3}$ and $C_{4}$.
A.3.3 The order-parameter correlation function